### v3

# Simplifying Hypergraph Neural Networks

Abstract
Hypergraphs are crucial for modeling higher-order interactions in real-world data. Hypergraph neural networks (HNNs) effectively utilise these structures by message passing to generate informative node features for various downstream tasks like node classification. However, the message passing block in existing HNNs typically requires a computationally intensive training process, which limits their practical use. To tackle this challenge, we propose an alternative approach by decoupling the usage of the hypergraph structural information from the model training stage. The proposed model, simplified hypergraph neural network (SHNN), contains a training-free message-passing block that can be precomputed before the training of SHNN, thereby reducing the computational burden. We theoretically support the efficiency and effectiveness of SHNN by showing that: 1) It is more training-efficient compared to existing HNNs; 2) It utilises as much information as existing HNNs for node feature generation; and 3) It is robust against the oversmoothing issue while using long-range interactions. Experiments based on six real-world hypergraph benchmarks in node classification and hyperlink prediction present that, compared to state-of-the-art HNNs, SHNN shows both competitive performance and superior training efficiency. Specifically, on Cora-CA, SHNN achieves the highest node classification accuracy with just $2\%$ training time of the best baseline.
Machine Learning, ICML
1 Introduction
Higher-order interactions involving more than two entities exist in various domains, such as co-authorships in social science (Han et al., 2009) and the spreading phenomena in epidemiology (Jhun, 2021). Hypergraphs, extending traditional graphs, effectively model these complex interactions with nodes representing entities and hyperedges connecting multiple entities (Bick et al., 2023). To leverage such structures for solving downstream tasks, the design of learning algorithms on hypergraphs has garnered increasing attention.
Inspired by the success of graph neural networks (GNNs) (Wu et al., 2020), most current research focuses on developing hypergraph neural networks (HNNs) with a message passing block (MP-Block) and a task-specific block [^1] For instance, in node classification, the task-specific block is implemented as a classifier, whereas in hyperlink prediction, it is designed as a predictor. Following previous works (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023), we focused on designing MP-Blocks that are compatible with various task-specific blocks.. The MP-Block enables information exchange between connected nodes to generate informative node features for the task-specific block (Feng et al., 2019; Bai et al., 2021; Chien et al., 2022; Wang et al., 2023a; Huang & Yang, 2021; Wang et al., 2023b; Duta et al., 2023; Telyatnikov et al., 2023). However, similar to other message passing neural networks (Frasca et al., 2020; wu2022graph), training the MP-Block makes loss computation interdependent for connected nodes, resulting in a computationally intensive training process for existing HNNs. This limits their practical applications, especially in semi-supervised learning that utilises the training and test data simultaneously (Joachims et al., 1999; Chapelle et al., 2009).
To address this challenge, inspired by the literature on efficient and effective GNNs (Frasca et al., 2020; Wu et al., 2019), our key solution is to move the integration of the hypergraph structural information from the training stage to the data preprocessing phase. This leads to a novel model called simplified hypergraph neural network (SHNN). In this model, a task-specific block is paired with a simplified message passing block (SMP-Block) that is training-free and can be precomputed in the data preprocessing phase. We develop the SMP-Block by simplifying the MP-Block of four state-of-the-art HNNs (Huang & Yang, 2021; Chen et al., 2022; Chien et al., 2022; Wang et al., 2023a) in two steps. First, we make the selected MP-Blocks training-free by eliminating their learnable parameters. Next, we enhance inference efficiency by removing non-linear activation functions and converting an $L$-layer MP-Blocks into a single propagation step. Remarkably, this two-step simplification unifies the chosen MP-Blocks into a single formulation, despite their different underlying design philosophies. We refer to this formulation as SMP-Block. Further, we thoroughly analyse SHNN to demonstrate its efficiency and effectiveness theoretically. Generally, we show three key properties: 1) SHNN is more training-efficient compared to existing HNNs; See Figure 1 for the comparison between the training pipelines of the HNN and our SHNN in node classification; 2) SHNN can utilise as much information as existing HNNs for generating node features; and 3) SHNN is robust against the oversmoothness issue when taking into account the long-range information.
The contributions of this work are summarised as follows:
$\bullet$ We develop SHNN, an efficient and effective model for hypergraph-structured data. To our knowledge, SHNN is the first model to integrate the hypergraph structural information in the data preprocessing stage, significantly enhancing training efficiency.
$\bullet$ We theoretically support the effectiveness and efficiency of SHNN by showing that: 1) It leads to remarkably low training complexity when solving hypergraph-related downstream tasks; 2) It utilises the same amount of information as existing HNNs for generating node features; and 3) It is robust against the oversmoothing issue when utilising long-range information.
$\bullet$ We conduct extensive experiments in both node classification and hyperedge prediciton tasks to compare SHNN with nine state-of-the-art HNNs. The empirical results show that the proposed SHNN exhibits both the competitive performance and superior training efficiency.
Figure 1: Training pipeline of HNN vs. SHNN in node classification. Top row: HNN uses a hypergraph structure to generate node features by a learnable MP-Block, which are then used by a classifier, with the MP-Block and the classifier being trained together. For brevity, we omit the MLP in HNN for the input node features. Bottom row: SHNN consists solely of the classifier being trained for node classification and the SMP-Block, which can be precomputed prior to classifier training.
2 Notation
Hypergraph. Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ be a hypergraph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, =$\mathcal{E}=\{e_{1},e_{2},\cdots,e_{m}\}$ is the hyperedge set, and $\mathbf{H}\in\{0,1\}^{n\times m}$ is an incidence matrix in which $\mathbf{H}_{ik}=1$ means that $e_{j}$ contains node $v_{i}$ and $\mathbf{H}_{ik}=0$ otherwise. Define $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal matrix of node degrees and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}\in\mathbb{R}_{\geq 0}^{m\times m}$ as a diagonal matrix of hyperedge degrees, where $\mathbf{D}_{\mathcal{H}^{\mathcal{V}}_{ii}}$ and $\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}$ are the number of hyperedges with $v_{i}$ and the number of nodes in $e_{k}$, respectively. In the following discussions, we assume the hypergraph does not have isolated nodes or empty/duplicate hyperedges. Moreover, we define: 1) The distance between two nodes on a hypergraph is the number of hyperedges in the shortest path between them, e.g., the distance is one if they are directly connected by a hyperedge; and 2) The $k$-hop neighbours of a node $v_{i}$ are all nodes with a distance of $k$ or less.
Graph and clique expansion. Let $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ be a graph, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{n}\}$ is the node set, and $\mathbf{W}\in\mathbb{R}_{\geq 0}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$ in which $\mathbf{W}_{ij}>0$ means that $v_{i}$ and $v_{j}$ are connected and $\mathbf{W}_{ij}=0$ otherwise. We set $\mathbf{D}\in\mathbb{R}_{\geq 0}^{n\times n}$ as a diagonal node degree matrix for $\mathcal{G}$, where $\mathbf{D}_{ii}$ is the sum of the $i$-th row of $\mathbf{W}$. Moreover, we denote the graph Laplacian of $\mathcal{G}$ as $\mathbf{L}=\mathbf{D}-\mathbf{W}$. Given a hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, its clique expansion is defined as a graph $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$, where $\mathcal{V}$ remains unchanged, and $\mathbf{W}_{ij}>0$ if and only if $v_{i}$ and $v_{j}$ are connected by a hyperedge on $\mathcal{H}$ and $\mathbf{W}_{ij}=0$ otherwise. Therefore, each hyperedge in $\mathcal{H}$ is a clique in $\mathcal{G}$.
Other notations. We set node features as $\mathbf{X}_{\mathcal{V}}\!=\![\mathbf{x}^{\top}_{v_{1}},\!\mathbf{x}^{\top}_{v_{2}},\!\cdots,\!\mathbf{x}^{\top}_{v_{n}}]^{\top}\in\mathbb{R}^{n\times d}$, which is a matrix that contains $d$-dimensional features. We denote functions or variables at the $l$-th layer of a model using the superscript $(l)$ and use $\oplus$ for concatenation. We use $\mathbf{\Theta}$ to represent the learnable weight matrix in a model, $\operatorname{MLP}(\cdot)$ for a multilayer perceptron (MLP), and $\sigma(\cdot)$ for non-linear functions (e.g., ReLU). In subsequent sections, we presume that the node features are $d$-dimensional unless otherwise noted.
3 Methodology
In this section, we aim to develop an efficient and effective model for hypergraph-structured data. In Subsection 3.1, we present the overall design of our simplified hypergraph neural network. Subsection 3.2 elaborates on the development of the simplified message passing block.
3.1 Overview
General formulation. Inspired by a previous work named SIGN (Frasca et al., 2020) in the GNN literature, we design a model, named simplified hypergraph neural network (SHNN), that can integrate the information about hypergraph structures into the data preprocessing stage, thus also removing the reliance on it at training. Mathematically, the SHNN can be summarised as the following formulation:
$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=S(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (1)
where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific block, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent [^2] For instance, in the node classification task, $\hat{\mathbf{Y}}\in\mathbb{R}^{n\times c}$ contains the logits for $c$ categories; and in the hyperlink prediction task, $\hat{\mathbf{Y}}\in\mathbb{R}^{m_{p}}$ contains the probability of $m_{p}$ potential hyperedges., $S(\cdot)$ denotes an MP-Block that contains only pre-defined parameters, allowing it to be precomputed before the training phase, and $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by $S(\cdot)$. In SIGN, $S(\cdot)$ is implemented by the combination of some graph filters. However, we cannot directly use existing hypergraph filters (Zhang et al., 2019; Qu et al., 2022) for two main reasons. Firstly, existing hypergraph filters are primarily designed for $k$-uniform hypergraphs, where all hyperedges have size $k$. Secondly, these filters require the use of an incidence tensor, which leads to prohibitively high memory usage for practical applications. Specifically, for a hypergraph with $n$ nodes and hyperedges of uniform size $k$, the corresponding incidence tensor has $n^{k}$ dimensions. Hence, instead of using existing methods, we draw inspiration from SGC (Wu et al., 2019) in the GNN literature to design a novel $S(\cdot)$ for hypergraphs, named simplified message passing block (SMP-Block), by simplifying MP-Blocks in state-of-the-art HNNs.
3.2 Simplified Message Passing Block
In this section, we first revisit existing hypergraph neural networks (HNNs), summarising these models and their MP-Blocks into general formulations. Furthermore, we develop the simplified message passing block (SMP-Block) by removing the learnable parameters and the non-linear function from the MP-Block in four state-of-the-art HNNs.
3.2.1 Revisiting Hypergraph Neural Networks
Table 1: Overview for four state-of-the-art HNNs and our SHNN. In this table, $\gamma_{U},\gamma_{E},\gamma_{D},\gamma^{\prime}_{l}\!\in\!(0,1)$ are hyperparameters and $\mathbf{I}\in\mathbb{R}^{d\times d}$ denotes an identity matrix. Moreover, $M$ is the training computational complexity of the task-specific block, $n$ is the node count, $m$ is the hyperedge count, $m^{\prime}$ is the number of edges in the clique expansion, $\|\mathbf{H}\|_{0}$ is the number of non-zero values in $\mathbf{H}$, $T$ is the number of training epochs, $L$ is the number of layers, and $d$ is the feature dimension.
Name Type Hypergraph-Wise MP Function Training Computational Complexity
UniGCNII (Huang & Yang, 2021) Direct MP $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{U}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\mathbf{\Theta}^{(l)}\Bigg{)}$ $\mathcal{O}(M+TL(n+m+\|\mathbf{H}\|_{0})d+TLnd^{2})$
Deep-HGNN (Chen et al., 2022) Direct MP $\mathbf{X}_{\mathcal{V}}^{(l)}\!=\!\sigma\Bigg{(}\Big{(}(\!1\!-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)}\Big{(}(1-\gamma^{\prime}_{D})\mathbf{I}+\gamma^{\prime}_{D}\mathbf{\Theta}^{(l)}\Big{)}\Bigg{)}$ $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$
AllDeepSets (Chien et al., 2022) Indirect MP $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
ED-HNN (Wang et al., 2023a) Indirect MP $\mathbf{X}_{\mathcal{V}}^{(l)}=\operatorname{MLP}\Bigg{[}(1-\gamma_{E})\operatorname{MLP}\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\operatorname{MLP}\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\operatorname{MLP}(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]}$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
SHNN (Ours) Direct MP $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$ $\mathcal{O}(M)$
General formulation. An HNN for downstream tasks is formulated as the following equation:
$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\hat{\mathbf{X}}_{\mathcal{V}}),~{}~{}\hat{\mathbf{X}}_{\mathcal{V}}=\Phi_{\mathbf{\Theta}}(\mathbf{X}_{\mathcal{V}},\mathbf{H}),$ (2)
where $\varphi_{\mathbf{\Theta}}(\cdot)$ is a task-specific block, $\hat{\mathbf{Y}}$ denotes the task-specific output whose dimension is task-dependent, $\Phi_{\mathbf{\Theta}}(\cdot)$ is a learnable message passing block, and $\hat{\mathbf{X}}_{\mathcal{V}}\in\mathbb{R}^{n\times d}$ denote the features generated by $\Phi_{\mathbf{\Theta}}(\cdot)$. The blocks $\varphi_{\mathbf{\Theta}}(\cdot)$ and $\Phi_{\mathbf{\Theta}}(\cdot)$ are trained together with the supervision given by the downstream tasks. We detail the message passing block in the next paragraph.
Message passing block. Generally, existing HNNs are based on the message passing mechanism, adapted from the graph machine learning literature (Xu et al., 2019; Bronstein et al., 2021). This mechanism makes models iteratively update the features of each node by aggregating the features of its neighbours. Mathematically, the features of a specific node $v_{i}$ are updated by the following propagation operators:
$\mathbf{x}_{v_{i}}^{(0)}=\operatorname{MLP}(\mathbf{x}_{v_{i}}),~{}~{}\mathbf{x}_{v_{i}}^{(l)}=\phi_{\mathbf{\Theta}}(\mathbf{x}_{v_{i}}^{(0)},\mathbf{x}_{v_{i}}^{(l-1)},\oplus_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}\mathbf{x}_{v_{j}}^{(l-1)}),$ (3)
where $\phi_{\mathbf{\Theta}}(\cdot)$ denotes a learnable node-wise message passing function that facilitates the information exchange between neighbouring nodes, and $\mathcal{N}_{\mathcal{H}_{v_{i}}}$ is the set of neighbours of $v_{i}$ on the hypergraph. Specifically, $\phi_{\mathbf{\Theta}}(\cdot)$ can be categorized into two types: direct and indirect message passing functions. The direct approaches aggregate features of neighbouring nodes to that of the target node directly (Huang & Yang, 2021; Wang et al., 2023b; Chen et al., 2022). In the indirect approaches, the features of neighbouring nodes are aggregated to that of the target node via some virtual nodes (Chien et al., 2022; Wang et al., 2023a; Yang et al., 2022). Mathematically these approaches can be formulated as:
$\mathbf{x}_{v_{i}}^{(l)}$ $=f_{3}\Big{(}f_{0}(\mathbf{x}_{v_{i}}^{(0)})+f_{1}(\mathbf{x}_{v_{i}}^{(l-1)})+\sum_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}f_{2}(\mathbf{x}^{(l-1)}_{v_{j}})\Big{)},$ (4a)
$\mathbf{x}_{v_{i}}^{(l)}$ $=g_{2}\Big{(}g_{0}(\mathbf{x}_{v_{i}}^{(0)})+\mathbf{p}^{\top}g_{1}(\mathbf{x}_{v_{i}}^{(l-1)},\oplus_{v_{j}\in\mathcal{N}_{\mathcal{H}_{v_{i}}}}\mathbf{x}_{v_{j}}^{(l-1)})\Big{)},$ (4b)
where $f_{0}(\cdot)$, $f_{1}(\cdot)$, $f_{2}(\cdot)$, $f_{3}(\cdot)$, $g_{0}(\cdot)$, $g_{1}(\cdot)$, $g_{2}(\cdot)$ represent seven learnable functions, and $\mathbf{p}\in\mathbb{R}^{n_{v}}$ is a vector used to aggregate the features of virtual nodes generated by $g_{1}(\cdot)$. Here, $n_{v}$ is a hyperparameter corresponding to the number of the virtual nodes, and Eq. (4a) and Eq. (4b) represent the direct and indirect approaches respectively. Further, we summarise four state-of-the-art HNNs in Table 1, and provide a more detailed review of existing HNNs in Appendix. In HNNs, an MP-Block typically consists of a combination of $L$ message-passing functions. In the following subsection, we introduce our SMP-Block by simplifying the MP-Blocks of HNNs in Table 1.
3.2.2 Simplifying Hypergraph Neural Networks
Our goal is to introduce the SMP-Block by simplifying the state-of-the-art HNNs presented in Table 1.
Learnable parameters elimination. To develop a training-free MP-Block, we first eliminate the learnable parameters from the message passing functions of the selected HNNs. Specifically, we replace the learnable matrices in these functions with identity matrices. This simplification allows us to represent the simplified message passing functions as follows:
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Big{(}(1-\gamma_{U})\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\tilde{\mathbf{D}}_{\mathcal{H}^{\mathcal{E}}}^{-1/2}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{X}_{\mathcal{V}}^{(l-1)}+\gamma_{U}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)},$ (5a)
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Big{(}(1-\!\gamma_{D}\!)\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{H}\!\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1/2}\mathbf{X}_{\mathcal{V}}^{(l-1)}\!+\!\gamma_{D}\mathbf{X}_{\mathcal{V}}^{(0)}\Big{)},$ (5b)
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Bigg{(}\mathbf{D}_{\mathcal{H}^{\mathcal{V}}}^{-1}\mathbf{H}\sigma\Big{(}\mathbf{D}_{\mathcal{H}^{\mathcal{E}}}^{-1}\mathbf{H}^{\top}\sigma(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)},$ (5c)
$\mathbf{X}_{\mathcal{V}}^{(l)}$ $=\sigma\Bigg{[}(1-\gamma_{E})\sigma\Bigg{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{V}}}\mathbf{H}\sigma\Big{(}\mathbf{D}^{-1}_{\mathcal{H}^{\mathcal{E}}}\mathbf{H}^{\top}\sigma(\mathbf{X}_{\mathcal{V}}^{(l-1)})\Big{)}\Bigg{)}+\gamma_{E}\mathbf{X}_{\mathcal{V}}^{(0)}\Bigg{]},$ (5d)
where Eq. (5a), Eq. (5b), Eq. (5c) and Eq. (5d) represent the simplified formulation of UniGCNII, Deep-HGNN, AllDeepSets and ED-HNN, respectively. Notably, $\sigma(\cdot)$ only denotes a general non-linear function without parameters, not implying that these simplified models use the same function.
Linearisation. Recently, some works in graph machine learning (Wu et al., 2019; Wang & Zhang, 2022) show that the message passing mechanism can operate more efficiently without non-linear activation functions while maintaining high effectiveness. Inspired by these works, we further remove the non-linearity from the functions represented by Eqs. (5$\sim$5d), thereby transforming an MP-Block with $L$ message passing functions into a single propagation step. For the sake of brevity, we present the simplification results as Proposition 3.1 and provide the mathematical proof for this Proposition in Appendix.
Proposition 3.1.
Let $\mathbf{X}_{\mathcal{V}}$ be input node features, $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of the MP-Block of an UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN with $L$ MP functions, and $\alpha\!\in\![\!0,1\!)$. Assume that learnable parameters and the non-linearity are removed from the block. For any hypergraph =$\mathcal{H}\!=\!\!\{\!\mathcal{V},\!\mathcal{E},\!\mathbf{H}\!\}$, there exists a clique expansion $\mathcal{G}\!=\!\{\!\mathcal{V},\mathbf{W}\!\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}\!=\!\Big{(}(1\!-\!\alpha)^{L}\!\mathbf{W}^{L}\!+\!\alpha\!\sum_{l=0}^{L-1}(1\!-\!\alpha)^{l}\!\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
According to Proposition 3.1, we design the formulation of our SMP-Block as the following equation:
$\hat{\mathbf{X}}=\mathbf{S}\mathbf{X}_{\mathcal{V}},$
where $\mathbf{S}=(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$. We define that the $L$ used to compute $\mathbf{S}$ is the number of layers of an SMP-Block. To intuitively understand the behaviour of an $L$-layer SMP-Block, we present the Lemma 3.1. For conciseness, we show the proof of this Lemma in Appendix.
Lemma 3.1.
Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\mathbf{H})$ be a hypergraph and $\mathcal{G}=(\mathcal{V},\mathbf{W})$ be its clique expansion with self-loops. Assume that $\alpha\in[0,1)$, and $\mathbf{S}\!=\!(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$. Then, for $i\neq j$, we have $\mathbf{S}_{ij}>0$ if and only if $v_{i}$ is an $L$-hop neighbour of $v_{j}$ on $\mathcal{H}$, and $\mathbf{S}_{ij}=0$ otherwise.
By Lemma 3.1, an $L$-layer SMP-Block enables the information exchange between any node $v_{i}$ and its $L$-hop neighbours on the hypergraph. In the next paragraph, we detail the specific designs of $\mathbf{S}$.
Operator design. To use the SMP-Block in our SHNN, the key is designing $\mathbf{W}$ to generate $\mathbf{S}$. We achieve this using a hyperedge-size-based edge weight that is defined as:
$\mathbf{W}_{{H}_{ij}}=\sum_{k=1}^{m}\frac{\delta(v_{i},v_{j},e_{k})}{\mathbf{D}_{\mathcal{H}^{\mathcal{E}}_{kk}}},$ (6)
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connects $v_{i}$ and $v_{j}$ and returns 0 otherwise. The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label. Consequently, this edge weight design can make the weighted clique expansion fit the homophily assumption, which is that connected nodes tend to be similar to each other (McPherson et al., 2001). We theoretically elaborate on this intuition in Appendix. Furthermore, inspired by previous works in the graph domain (Wu et al., 2020), we generate $\mathbf{S}$ based on a symmetrically normalised $\mathbf{W}_{H}$ with the self-loop. Specifically, we generate the $\mathbf{S}$ by:
$\mathbf{S}=(1-\alpha)^{L}(\tilde{\mathbf{D}}_{H}^{-1/2}\tilde{\mathbf{W}}_{H}\tilde{\mathbf{D}}_{H}^{-1/2})^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}(\tilde{\mathbf{D}}_{H}^{-1/2}\tilde{\mathbf{W}}_{H}\tilde{\mathbf{D}}_{H}^{-1/2})^{l},$ (7)
where $\tilde{\mathbf{W}}_{H}=\mathbf{W}_{H}+\mathbf{I}_{n}$ and $\tilde{\mathbf{D}}_{H}\in\mathbb{R}^{n\times n}$ is the diagonal node degree matrix of $\tilde{\mathbf{W}}_{H}$. In the next section, we present the theoretical analysis to support the efficiency and effectiveness of our SHNN.
4 Analysis
In this section, we begin by showing the efficiency of the proposed SHNN by comparing its training complexity with the HNNs. Additionally, we demonstrate the effectiveness of our SHNN by analyzing its information utilisation and its robustness against the oversmoothing issue.
Training complexity in downstream tasks. As illustrated by Eq. (1), the SHNN in downstream tasks requires training only the task-specific block. Additionally, since the SHNN relies on a pre-defined one-step propagation operator, its computations can be handled during data pre-processing. Consequently, the training complexity of the SHNN-based framework is determined solely by the task-specific block. In contrast, on the basis of Eq. (2), the HNN involve training both the HNN and the task-specific block. The HNN includes complex learnable operations that facilitate message passing between connected nodes on a hypergraph, inherently making the training of the HNN-based framework more complex than the SHNN. See Figure 1 for an example that qualitatively compares the training pipelines of the SHNN and HNN in node classification. To provide a quantitative comparison, we summarize the training computational complexity of the SHNN and the four state-of-the-art HNNs in Table 1. This table shows that the training complexity of the SHNN is consistently lower than that of the HNN, theoretically confirming the training efficiency of the proposed SHNN.
Information utilisation. The core strength of existing HNNs lies in their ability to utilize the information embedded within hypergraph structures to generate powerful node features for downstream tasks. Consequently, we analyse the effectiveness of our proposed SHNN by comparing its information utilization capabilities with those of existing HNNs and conclude that our methods utilise the same amount of information as exising HNNs without any loss, which is summarised in Proposition 4.1.
Proposition 4.1.
Let $H_{0}^{v^{L}_{i}}$ be the entropy of information used by an HNN with $L$ message passing functions in generating the features of a node $v_{i}$, $H_{1}^{v^{L}_{i}}$ be the entropy of information used by an SHNN with an $L$-layer SMP-Block for the same purpose, and $H_{2}^{v^{L}_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours on the hypergraph. Then, $H_{0}^{v^{L}_{i}}=H_{1}^{v^{L}_{i}}=H_{2}^{v^{L}_{i}}$
This proposition theoretically demonstrates that SHNN is as effective as existing HNNs in utilizing information from pre-defined hypergraphs to generate node features. Specifically, both SHNN and existing HNNs achieve this by aggregating neighbourhood information.
Robustness against the oversmoothing issue. According to Lemma 3.1, SHNN can leverage global interactions within the given hypergraph to generate node features by deepening the SMP-Block. However, as highlighted in a recent study (Chen et al., 2022), the message-passing-based models on hypergraphs may suffer from the oversmoothing issue. This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which consequently degrades the performance of the learning framework in downstream tasks. Consequently, to further support the effectiveness of SHNN, we theoretically analyse its robustness against the oversmoothing issue. Inspired by previous works in the graph machine learning literature (Yang et al., 2021; zhang2020revisiting), we demonstrate the following Proposition:
Proposition 4.2.
Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ denote a hypergraph, $\mathcal{G}=\{\mathcal{V},\mathbf{W}_{H}\}$ be its clique expansion with edge weights computed by Eq. (6), $\mathbf{L}$ be the graph Laplacian matrix of $\mathcal{G}$ computed by a symmetrically normalised and self-loops added $\mathbf{W}_{H}$, $\mathbf{X}_{\mathcal{V}}$ represent the input node features, and $\alpha\in[0,1)$ be a hyperparameter. We define that $F(\mathbf{X}^{\prime})=\operatorname{tr}(\mathbf{X}^{\prime\top}\mathbf{L}\mathbf{X}^{\prime})+\frac{\alpha}{1-\alpha}\operatorname{tr}[(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})^{\top}(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})]$, and $F_{min}$ as the global minimal value of $F(\mathbf{X}^{\prime})$ for $\mathbf{X}\in\mathbb{R}^{n\times d}$. Assume that: 1) $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$; 2) $\mathbf{S}$ is computed by Eq. (7); 3) $\alpha>0$; and 4) $L\to+\infty$. Then, $F(\hat{\mathbf{X}}_{\mathcal{V}})=F_{min}$.
Notably, minimising the first term of $F(\mathbf{X}^{\prime})$ enhances feature similarity among connected nodes on $\mathcal{H}$, and minimising the second term of $F(\mathbf{X}^{\prime})$ encourages the generated features of each node to retain the distinct information from its input features. Therefore, node features that can minimise $F(\mathbf{X}^{\prime})$ capture two key properties: 1) neighbouring nodes on $\mathcal{H}$ have similar features; and 2) each node contains unique information reflecting its individual input characteristics.
Based on Proposition 4.2, a deep SMP-Block tends to assign similar features to connected nodes on a hypergraph while maintaining each node’s unique input information. Hence, even with an extremely large number of layers, the SMP-Block can preserve the unique input information of each node, making it robust against the oversmoothing issue. This property further ensures the effectiveness of our SHNN. In Section 6, we empirically support the analysis presented in this section.
5 Related Work
Efficient learning with hypergraphs. Recent efforts have risen to utilize existing hypergraphs to enhance node feature learning, thereby benefiting downstream tasks (Bick et al., 2023). The predominant models in this area are hypergraph neural networks (HNNs) with a message passing block that enables information exchange between connected nodes (Antelmi et al., 2023). Similar to their graph counterparts (Wu et al., 2019; Frasca et al., 2020), HNNs suffer from low learning efficiency. There are two recent works for this limitation (Feng et al., 2024; Tang et al., 2024). (Feng et al., 2024) distills an HNN named HGNN (Feng et al., 2019) into an MLP during training, while (Tang et al., 2024) integrates hypergraph structural information into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge. These works focus on reducing model inference complexity via a dedicated training process, suitable for scenarios where training and test data are clearly separated. However, many existing models are applied to semi-supervised hypergraph node classification settings that involve the simultaneous use of training and test data (Feng et al., 2019; Bai et al., 2021; Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a, b; Duta et al., 2023; Telyatnikov et al., 2023; Joachims et al., 1999; Chapelle et al., 2009). Therefore, reducing the training complexity of hypergraph-based machine learning models remains an important yet underexplored challenge. We address this challenge with a novel model named simplified hypergraph neural network (SHNN), which is the first to incorporate hypergraph structural information during data preprocessing. This integration significantly improves training efficiency.
Table 2: Dataset statistics. CE homophily is the homophily score based on clique expansions of hypergraphs; details of this score can be found in Appendix. Moreover, NC denotes the hypergraph node classification task and HP represents the hyperedge prediction task.
Name # Nodes # Hyperedges # Features # Classes CE homophily Node Definition Hyperedge Definition Task
Cora-CA 2708 1072 1433 7 0.803 Paper Co-authorship NC&HP
DBLP-CA 41302 22363 1425 6 0.869 Paper Co-authorship NC
Citeseer 3312 1079 3703 6 0.893 Paper Co-citation NC & HP
Congress 1718 83105 100 2 0.555 Congressperson Legislative-Bills-Sponsorship NC
House 1290 340 100 2 0.498 state-of-the-art Committee NC&HP
Senate 282 315 100 2 0.509 state-of-the-art Legislative-Bills-Sponsorship NC
6 Experiments
6.1 Experiment Setup
Task description. We conduct experiments in two downstream tasks: hypergraph node classification and hyperlink prediction. In hypergraph node classification (Wang et al., 2023a, b; Duta et al., 2023), we are given the hypergraph structure $\mathbf{H}$, node features $\mathbf{X}_{\mathcal{V}}$, and a set of labelled nodes with ground truth labels $\mathbf{Y}_{lab}=\{\mathbf{y}_{v}\}_{v\in\mathcal{V}_{lab}}$, where $\mathbf{y}_{v_{i}}\in\{0,1\}^{c}$ be a one-hot label. Our objective is to classify the unlabeled nodes within the given hypergraph. In hyperlink prediction (Chen & Liu, 2023), we are given node features $\mathbf{X}_{\mathcal{V}}$, a set of observed real and fake hyperedges $\mathcal{E}_{ob}$, and a set of potential hyperedges $\mathcal{E}_{p}$. The task requires the model to distinguish between real and fake hyperedges in $\mathcal{E}_{p}$ using $\mathbf{X}_{\mathcal{V}}$ and $\mathcal{E}_{ob}$.
Dataset and baseline. For the hypergraph node classification task, we conduct experiments on six real-world hypergraphs: Cora-CA, DBLP-CA, Citeseer, Congress, House, and Senate. We use nine HNNs as baselines: HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), HNHN (Dong et al., 2020), UniGCNII (Huang & Yang, 2021), AllDeepSets (Chien et al., 2022), AllSetTransformer (Chien et al., 2022), PhenomNN (Wang et al., 2023b), ED-HNN (Wang et al., 2023a), and Deep-HGNN (Chen et al., 2022). For the hyperlink prediction task, we compare our SHNN with AllSetTransformer, PhenomNN, ED-HNN, and Deep-HGNN on three real-world hypergraphs: Cora-CA, Citeseer, and House. The dataset statistics are summarized in Table 2, with detailed descriptions of datasets available in Appendix.
Metric. Following previous works (Wang et al., 2023a; Chen & Liu, 2023), we evaluate models for node classification and hyperlink prediction with classification accuracy and the Area Under the ROC Curve (AUC), respectively. We evaluate the efficiency of the models by the relative training time $r_{f}$, defined as $r_{f}=t_{f}/t_{S}$, where $t_{f}$ is the training time to achieve optimal performance for the evaluated model, and $t_{s}$ is the training time for optimal performance of our SHNN.
Implementation. For node classification, we follow previous works (Wang et al., 2023a; Duta et al., 2023) to use a 50%/25%/25% train/validation/test data split and adapt the baseline classification accuracy from them [^3] Since there are no reported results for Deep-HGNN under the chosen data split, we use our own results.. Additionally, similar to these works, we implement the classifier based on MLP for our SHNN and report the results from ten runs. For hyperlink prediction, existing works (Chen & Liu, 2023) primarily focus on the design of the prediction head, with no reported results for applying our baseline HNNs to this task; so we report all results conducted by ourselves. We employ the deep set function implemented by (Chien et al., 2022) as the prediction head for our method and baseline methods due to its simplicity. We use a 50%/25%/25% train/validation/test data split and ensure each split contains five times as many fake hyperedges as real hyperedges. Both SHNN and HNNs only use real hyperedges in training and validation sets to do message passing. All experiments were performed on an RTX 3090 with PyTorch, and our code with running commands is at https://anonymous.4open.science/r/SHNN. More details are in Appendix.
6.2 Comparison with Baselines
Table 3: Node classification accuracy (%) for HNNs and SHNN. The best result on each dataset is highlighted in bold font. The second and third highest accuracies are marked with an underline.
Cora-CA DBLP-CA Citeseer Congress House Senate Avg. Mean
HGNN 82.64 ± 1.65 91.03 ± 0.20 72.45 ± 1.16 91.26 ± 1.15 61.39 ± 2.96 48.59 ± 4.52 74.56
HCHA 82.55 ± 0.97 90.92 ± 0.22 72.42 ± 1.42 90.43 ± 1.20 61.36 ± 2.53 48.62 ± 4.41 74.38
HNHN 77.19 ± 1.49 86.78 ± 0.29 72.64 ± 1.57 53.35 ± 1.45 67.80 ± 2.59 50.93 ± 6.33 68.12
UniGCNII 83.60 ± 1.14 91.69 ± 0.19 73.05 ± 2.21 94.81 ± 0.81 67.25 ± 2.57 49.30 ± 4.25 76.62
AllDeepSets 81.97 ± 1.50 91.27 ± 0.27 70.83 ± 1.63 91.80 ± 1.53 67.82 ± 2.40 48.17 ± 5.67 75.31
AllSetTransformer 83.63 ± 1.47 91.53 ± 0.23 73.08 ± 1.20 92.16 ± 1.05 51.83 ± 5.22 69.33 ± 2.20 76.93
PhenomNN 85.81 ± 0.90 91.91 ± 0.21 75.10 ± 1.59 88.24 ± 1.47 70.71 ± 2.35 67.70 ± 5.24 79.91
ED-HNN 83.97 ± 1.55 91.90 ± 0.19 73.70 ± 1.38 95.00 ± 0.99 72.45 ± 2.28 64.79 ± 5.14 80.30
Deep-HGNN 84.89 ± 0.88 91.76 ± 0.28 74.07 ± 1.64 93.91 ± 1.18 75.26 ± 1.76 68.39 ± 4.79 81.38
SHNN (ours) 86.54 ± 1.32 91.80 ± 0.30 74.82 ± 1.67 95.09 ± 0.89 76.29 ± 1.99 70.42 ± 2.74 82.50
Figure 2: The relative training time required for HNNs and SHNN to achieve optimal performance.
(a) Cora-CA.
(b) Citeseer.
(c) House.
Figure 3: Hyperlink prediction AUC (%) and relative training time for HNNs and SHNN.
Node classification. We summarize the classification accuracy and relative training time of the SHNN and HNNs in Table 3 and Figure 2, respectively. Table 3 demonstrates that SHNN not only leads to the best results across four datasets (Cora-CA, Congress, House, and Senate) but also results in the highest average mean accuracy overall. These findings confirm the effectiveness of SHNN in generating powerful node features for node classification. As discussed in Section 4, we attribute this effectiveness to SHNN’s ability to match the power of existing HNNs in utilizing neighbourhood information to generate node features, while requiring the optimization of fewer parameters. This makes the SHNN more efficient in utilizing training data and reduces the risk of overfitting, thereby outperforming more complex HNN counterparts. On the other hand, as shown in Figure 2, the SHNN consistently needs less training time to achieve superior performance compared to the HNNs. This underscores the high training efficiency of SHNN in downstream tasks.
Hyperlink prediction. The results for the hyperlink prediction task are summarized in Figure 3. These results exhibit a pattern similar to that observed in the hypergraph node classification task: the SHNN framework outperforms the HNNs while requiring less training time. An interesting observation is that the training efficiency improvement brought by SHNN is smaller in the hyperlink prediction task compared to the node classification task. The intuition behind this observation is that the task-specific block used for hyperlink prediction is more computationally expensive than the one used for hypergraph node classification. Specifically, the hyperlink predictor needs to generate a prediction for each hyperedge by aggregating the features of the nodes connected by the hyperedge, whereas the node classifier only needs to perform forward propagation for each individual node. We theoretically discuss that the efficiency improvement provided by SHNN is inversely correlated with the complexity of the task-specific block in Appendix.
6.3 Ablation study
We conduct a series of ablation studies on both the SPM-Block and the general SHNN model in node classification on the Cora-CA, Citeseer, House, and Senate datasets.
(a) Cora-CA.
(b) Citeseer.
(c) House.
(d) Senate.
Figure 4: The ablation study on the hyperparameters of SPM-Block in node classification.
Table 4: The ablation study on the components of SHNN in node classification. The best result on each dataset is highlighted in bold font.
SMP-Block Weighted $\mathbf{S}$ Cora-CA Citeseer Senate House Avg. Mean
$\bm{\times}$$\bm{\times}$ 77.28 ± 1.57 73.89 ± 1.75 68.31 ± 3.40 74.43 ± 2.44 73.48
$\bm{\surd}$$\bm{\times}$ 84.90 ± 1.55 74.75 ± 1.62 68.87 ± 4.57 75.57 ± 1.81 76.02
$\bm{\surd}$ $\bm{\surd}$ 86.54 ± 1.32 74.82 ± 1.67 70.42 ± 2.74 76.29 ± 1.99 77.02
Ablation study on the hyperparameters of SPM-Block. We focus on two key hyperparameters of the SPM-Block: the number of layers and the $\alpha$. As illustrated in Figure 4, models with $\alpha>0$ exhibit more consistent accuracy as the number of layers increases compared to models with $\alpha=0$. We attribute this consistency to the robustness discussed in Proposition 4.2, where a positive $\alpha$ helps retain the distinct information of each node, thereby mitigating the oversmoothing issue. Additionally, we observe that a positive $\alpha$ significantly enhances performance in the House and Senate datasets, which are more heterophilic, compared to the Cora-CA and Citeseer datasets. This suggests that retaining the distinct information of each node is more beneficial for heterophilic datasets.
Ablation study on the components of SHNN. In this study, we examine the influence of the SMP-Block and the hyperedge-size-based clique expansion on the performance of the SHNN in node classification. Without the SMP-Block, the model functions as a pure MLP. We define $\mathbf{S}$ based on hyperedge-size-based clique expansion as weighted $\mathbf{S}$; in contrast, for the framework without weighted $\mathbf{S}$, $\mathbf{S}$ is generated using a $\mathbf{W}$ with all positive values set to one. As shown in Table 4, the inclusion of SMP-Block improves classification accuracy across all datasets, demonstrating that SMP-Block effectively incorporates hypergraph structural information into node features. Furthermore, adding weighted $\mathbf{S}$ further enhances performance, highlighting the informativeness of the hyperedge-size-based clique expansion. Thus, the components we designed significantly aid SHNN in leveraging useful information from the given hypergraph structure.
7 Conclusion & Limitation
In this paper, we propose a novel model called the simplified hypergraph neural network (SHNN). The key innovation of SHNN is an original training-free messgea passing block tailored for data on hypergraphs. We present both theoretical and empirical results demonstrating that SHNN can solve hypergraph-related downstream tasks with remarkable efficiency and effectiveness. To our knowledge, SHNN is the first model to integrate the hypergraph structural information in the data preprocessing stage, which significantly enhances training efficiency. The proposed SHNN can advance the field of hypergraph machine learning research by not only improving the efficiency of using observed hypergraph structures to solve downstream tasks, but also serving as a simple starting point for the development of future hypergraph machine learning models.
Limitation and future work. Learning tasks on hypergraphs can generally be defined at the node, hyperedge, and hypergraph levels. In this paper, we empirically test our SHNN on node-level and hyperedge-level tasks, with the main limitation being the lack of empirical results on hypergraph-level tasks. To our knowledge, designing task-specific blocks to generate hypergraph-level features based on existing node features itself remains a challenging topic in the literature. Therefore, we consider the application of our SHNN to hypergraph-level tasks as future work.